OpenStudy (anonymous):

Graph the linear system and determine how many solutions it has: no solution, exactly one solution or infinitely many solutions x+2y=-2 -3x-6y=6 please help me determine the ordered pairs..

OpenStudy (anonymous):

We have infinitely way to answer !!!:)

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

Look : One of that ways !: we have : x+2y=-2 -3x-6y=6 and can say : 3x+6y=-6 -3x-6y=6 Sum them : 0+0=0 So what do you think ? We can get it with many ways !!!:)

OpenStudy (anonymous):

how do i graph it?

OpenStudy (anonymous):

What ??? we cant !!! We dont have any answer !!!

OpenStudy (anonymous):